Accounting

Kane Biotech was preparing the annual financial statements and, as part of the year-end procedures, assessed the assets and prepared the following alphabetized schedule based on adjusted values at December 31, 2020: (Do not round intermediate calculations. Round final answers to nearest whole dollar.)

AssetDate of PurchaseDeprec. Method*CostResidual ValueUseful LifeAccum. Deprec.Recoverable Amount








EquipmentMay 1/15Units$63,000
$3,160
8,800 units$39,440
$6,160

FurnitureJun. 28/15DDB
30,000

3,800
8 yrs
25,169

5,650

LandApr. 5/15N/A
103,000

N/A
N/A
N/A

119,800

Office buildingApr. 5/15SL
113,000

33,200
15 yrs
39,200

70,500

WarehouseApr. 5/15SL
82,000

28,000
20 yrs
30,938

63,100


















*DDB = Double-declining-balance; SL = Straight-line; Units = Units-of-production; N/A = Not applicable

Required:
1. 
Record any impairment losses at December 31, 2020. Assume Kane Biotech has recorded no impairment losses in previous years.





2. Record depreciation for each asset at December 31, 2021. Assume that there was no change in the residual values or useful lives regardless of any impairment losses that might have occurred. The equipment produced 2,700 units during 2021.


please do all calculations on page (journal entries)
Answer #1

1. Record any impairment losses at December 31, 2020. Assume Kane Biotech has recorded no impairment losses in previous years.

Asset Cost Accumulated Depreciation Carrying Value Recoverable Amount Impairment Revised Carrying Amount
[A] [B] [C = A-B] [D] [E = C-D]
Equipment $63,000 $39,440 $23,560 $6,160 $17,400 $6,160
Furniture $30,000 $25,169 $4,831 $5,650 NA $4,831
Land $103,000 NA NA $119,800 NA $119,800
Office Buildings $113,000 $39,200 $73,800 $70,500 $3,300 $70,500
Warehouse $82,000 $30,938 $51,062 $63,100 NA $51,062

Prepare entry as follows:

Account Titles Debit Credit
Impairment Loss $20,700
Equipment $17,400
Office Buildings $3,300

____________________________________________________________

2. Record depreciation for each asset at December 31, 2021. Assume that there was no change in the residual values or useful lives regardless of any impairment losses that might have occurred.

Working as follows:

Equipment [Carrying Value as on 31-3-2020] Units Produced Total Units [56160 x 2700 units] 8800 units =$1890 Furniture=Carryi

Carrying value-Residual value Office Building= Remaining life $70500-$33200 10 years = $3730 Carrying value-Residual value Wa

Account Titles Debit Credit
Impairment Loss $20,700
Equipment $17,400
Office Buildings $3,300
Depreciation $8,365.20
Equipment $1,890
Furniture $1,207.75
Office Building $3,730
Warehouse $1,537.45